# Problem Solving Through Human-AI Preference-Based Cooperation

### v2

@@ 3.2 Latent utility
We assume that the construction space is equipped with a latent utility function reflecting the preferences of the expert, i.e., the quality of solutions as perceived by the expert. In general, “quality” may refer to various dimensions or criteria, and different objectives might be pursued at the same time; we formalize this with a multi-dimensional utility function $u({\bm{x}})=(u_{1}({\bm{x}}),\ldots,u_{K}({\bm{x}}))^{\top}$ comprised of local utility functions $u_{i}$. For example, a computer program could be rated by average runtime or memory consumption. The local utility functions can be combined into a scalar utility function $U:\,{\mathcal{X}}\rightarrow\mathbb{R}$ via a suitable aggregation operator.
Various factors influencing the quality of candidate solutions can be distinguished, notably hard and soft constraints. Hard constraints refer to (functional) properties that qualify a candidate as a valid solution. For example, a computer program should properly compile and not contain any syntax errors. Even if invalid solutions should normally be considered useless, the abstract notion of utility is flexible enough to distinguish different levels of invalidity. For example, a non-executable computer program may still have a non-zero utility if the error can easily be fixed by the expert. In any case, hard constraints will normally not identify a solution in a unique way. For example, there are many computer programs that are functionally equivalent in the sense of having the same input-output behavior. Soft constraints refer to criteria that make a solution more or less desirable such as the length of a computer program and its time and memory consumption.
In general, the utility (be it in the form of the multi-dimensional utility function $u$ or the scalar utility function $U$) is not known to the AI, nor is the expert explicitly aware of it. Rather, this utility is latent and underlies the expert’s preference feedback. From this preference feedback, the AI can learn an approximation $\hat{U}$. The AI’s goal is then to construct a solution ${\bm{x}}^{*}$ that maximizes $\hat{U}$, or which is at least close to the maximizer, while simultaneously improving the approximation quality of $\hat{U}$. The utility $U$ also induces utilities on higher levels of abstraction. For example, one way to “lift” a utility function from level ${\mathcal{X}}_{j}$ to the more abstract level ${\mathcal{X}}_{j-1}$ is via aggregation: $U({\bm{x}}^{\prime})=\alpha(\{U({\bm{x}})\,|\,{\bm{x}}\in{\mathcal{X}}_{j},f_{j}({\bm{x}})={\bm{x}}^{\prime}\})$, where $\alpha$ is an appropriate aggregation function (Grabisch et al., 2009).
@@ 3.3 Interaction and preference-based search
To guide the search, human and AI can communicate via natural language; e.g., the AI agent may ask the expert for feedback or explicit advice. Alternatively, the expert may actively intervene, for example by critiquing or modifying a candidate solution. A third type of interaction, particularly important in the context of HAI-Co2, is driven through preferential feedback: By informing the AI agent about the quality of candidate solutions, the expert provides hints at presumably more promising (and, likewise, less promising) regions of the construction space, and hence suggests promising “search directions” to the AI agent. To give an example from our case study (Section 5), the expert compares two competing candidate solutions with each other (e.g., whether a modification has improved a solution or made it worse) and provides this feedback (in natural language) to the AI agent for the next iteration. The AI agent utilizes the feedback to improve its approximation $\hat{U}$ of the latent utility function, which is an important element of its informational state.
The way in which the AI agent and the human expert cooperate with each other is defined in the form of a protocol. Among other things, the protocol clarifies the type of queries and responses on the two sides (AI agent and human expert) and the (preference) feedback that can be given by the expert.
@@
HAI-Co2 comprises a broad variety of human-AI interaction in natural language as well as categorical choices and active modification of the candidate solution by the human expert. The search policy $\pi$ is designed to generate a (locally) optimal candidate solution based on the immediate as well as historical feedback, thereby adapting to the preference signals from the human expert user. The hierarchical abstraction of the search space facilitates a modular modification of the complex candidate solution. As we will see in the case study (Section 5), HAI-Co2 also allows for incorporating creative components into the generation of candidate solutions, for example through the injection of randomness in the heuristic search process or the refinement of abstract into more concrete solutions.
@@ 4 Challenges and practical issues
Specification of abstraction hierarchy. Wwo core components of HAI-Co2 are (i) the abstraction hierarchy of the construction space and (ii) the neighborhood structure that facilitates preference-based search. A synergistic implementation of (i) and (ii) poses a non-trivial research challenge. In the domain of code generation, Le et al. (2024) propose a modular code generation approach to circumvent this challenge: they generate a chain-of-thought style intermediate description of the subtasks followed by modular codes implementing each of them. Such a hierarchical generation approach can be extended to generalized solution co-construction. However, relying on purely natural language-based intermediate representations limits the utility of the hierarchy. The choice of abstraction representation is domain-specific: UML descriptions are a suitable abstraction for code generation, but not for an AI scientific assistant. The abstraction specification must adhere to the neighborhood structure on all levels of abstraction; e.g., if two points are neighbors, their abstractions must also be neighbors. This poses additional constraints on the choice of abstraction, the choice of neighborhood structure, and the choice of refinements between two given levels of abstraction. Ideally, all levels of abstraction must be suitable for human preference expression: an “unnatural” choice of abstraction can render the co-construction tiresome for the human expert, negatively affecting productivity.
@@
Specification of informational state. HAI-Co2 utilizes an informational state to keep track of relevant information in the interaction history. Given that the search policy is conditioned on it, an efficient representation of the informational state is a core challenge of HAI-Co2. Typically such interactive co-constructions are expected to span a long sequence context. While we have observed a significant surge in the context-size of present-day generative AI (e.g., GPT-4 Turbo can handle up to 128K tokens in the input prompt), recent research has questioned the effective usability of such very long context information (Liu et al., 2024). The representation of the informational state needs to be compatible with the abstraction specification of the construction space as well as well as the choice of how preference signals from the human expert are encoded. This is particularly important as the reflection of any preference signal upon the candidate solution is manifested via the informational state – an unreliable update of the informational state subsequently worsens the solution quality and may result in a divergent search.
@@
Communication in natural language. When humans co-construct a solution, communication in natural language plays an important role. Natural language is a powerful and at the same time succinct medium for conveying information. Given the expressivity of natural language, human and AI agent can easily communicate different options of how to improve the current solution, both at a detailed level and in more abstract terms. Similarly, preference learning is facilitated by natural language since many preferences are easily specified in natural language. The challenge here is that the language capabilities of LLMs have advanced to an impressive level for the general domain, but this does not apply to complex expert domains (Magesh et al., 2024; Hager et al., 2024; Anand et al., 2024).
@@
Multimodal human-AI interaction. Natural language-based interaction is not the ideal channel for all types of preference. Categorical preference can be communicated more simply by pointing towards the better solution. Thus, we would like to incorporate multiple types of preference into HAI-Co2. This poses the challenge of aligning these multiple modes of feedback with each other. For example, the human expert may express the need for a security feature in a software engineering problem explicitly, or they can express it implicitly by choosing a candidate solution that includes the feature over one that does not. The AI needs to extract equivalent preference information in these two scenarios. Contemporary research in recommender systems that deal with modeling user preferences on multiple item modalities (Guo et al., 2018; Xu et al., 2021) can serve as a starting point. However, the relative complexity and nuances of preference in the case of HAI-Co2 hinder a trivial extension of recommendation-oriented solutions.
@@
Dynamic user preference. Current techniques of aligning neural AI systems to human preferences, broadly referred to as RLHF (Reinforcement Learning from Human Feedback), typically involve a two-stage process: learning a reward model on preference data followed by fine-tuning a foundation model (often an LLM or a diffusion model) upon reward supervision from the reward model (Kaufmann et al., 2024). This setup is fundamentally limited to static adaptation in the regime of expert-domain co-construction; a single model of human preference is imitated by the agent that cannot adapt to the personalized preference of the human expert. In practice, user preferences are dynamic and evolve over time and are, as discussed by Lichtenstein & Slovic (2006), commonly constructed during elicitation and as such highly influenced by the elicitation process. This is a fundamental challenge in co-construction problems, where the AI agent must adapt to the evolving preference of the human expert. Multi-turn RLHF (Zhou et al., 2024), though it extends the context of preference-adherence to an iterative, conversational regime, cannot solve the challenge of dynamically evolving user preference. The PAL framework, proposed by Chen et al. (2024), provides a partial solution to our problem via personalized modeling of static human preferences. Unlike traditional policy learning, HAI-Co2 motivates a reward-free exploration of the solution space (Jin et al., 2020). In-context reinforcement learning can pave the way towards handling dynamic preference signals (Yang et al., 2024; Lee et al., 2024). However, the action space in the scope of HAI-Co2 overlaps with the generation of multiple hierarchical views of the candidate solution, rendering the problem much harder than existing work on in-context RL. Prior work with LLMs showcases the possibilities of using them as in-context agents, though exploration abilities will need fine-tuning-based interventions (Krishnamurthy et al., 2024).
@@
  • Quality of the solution should be evaluated using domain-specific measures; irrespective of the process of co-construction or human preference, the solution must fulfill some objective criteria of correctness.
@@
  • Complexity of co-construction includes the computational complexity of generation and preference-based search – resulting in potentially high computational cost – and the cognitive complexity of the framework – resulting in cognitive load for the expert user. The latter demands significant research efforts from a multi-disciplinary approach to ensure that automated assistants truly bring value to the expert.
@@
Given that human experts are costly and have limited time, LLM-based simulation of human-AI interaction may facilitate large-scale evaluation (Tamoyan et al., 2024). Even though many of our four evaluation criteria seem to demand human evaluation, we conjecture that the development of artificial critic models (McAleese et al., 2024), with human value alignment, will be an important research direction in the future.
Figure 3: Instantiation of HAI-Co2 for the problem of building a double pendulum simulation. A solution is co-constructed through human-AI cooperation as follows. 1 The specification module takes the initial problem statement as input and generates a natural language description of the software. 2 The UML generation module generates multiple candidate UML descriptions from the generated specification using stochastic decoding. 3 Pairs of candidate UMLs are provided to the Preference learning module, which chooses a winner for each pair and generates a justification of its choice. We simulate a tournament 4 using the preference learning module in this pairwise manner and end up with a pair of “best” UML candidates. 5 The code generation module independently generates two Python programs, one for each UML candidate. Upon execution of these two candidate programs, the expert user provides their feedback 6 by choosing one of them as a “better” candidate, along with (optional) justification behind their choice and (optional) additional requirements. The human feedback, along with a summary of the tournament choices made by the preference learning module, are then used by the UML generation module to generate a new set of candidate UMLs.
5 An exemplary implementation of HAI-Co2
@@
In this section, we provide a case study of HAI-Co2, tailored to code generation as a co-construction problem. This case study does not claim scientific rigor on its own; instead, we use it to demonstrate what prior findings (see Section 1 and Section 2) already establish.
@@
The initial problem description is underspecified. During the cooperation, the user can introduce new requirements, ask for modifications to the already generated code, and so on. We approximate different aspects of HAI-Co2  (the surjective mappings between different abstraction hierarchies of the construction space, policy and heuristic search strategy and the multi-dimensional utility function) using baseline implementation strategies for ease of demonstration. Future research endeavors should be directed to more in-depth implementation of these features.
@@
In this example, the construction space ${\mathcal{X}}$ consists of the set of all Python programs. Three distinct levels of abstraction are implemented: a specification of the simulation software in natural language (${\mathcal{X}}_{0}$), a UML description of the software (${\mathcal{X}}_{1}$), and finally, the Python program (${\mathcal{X}}_{2}$) itself. The abstraction refinements $f^{(-1)}_{1}$ and $f^{(-1)}_{2}$ (as introduced in Section 3) are implemented using suitably prompted instances of GPT-4 Turbo that we denote as UML generation module and Code generation module, respectively; while the former produces a (stochastic) set of refinements in UML given a natural language specification, the latter generates Python implementations of a given UML description. The informational state ${\mathcal{I}}$ is realized at different levels of abstraction within the contexts of these LLMs. We do not implement a concrete realization of the policy $\pi$; instead, we rely on the limited abilities of LLM instances to explore and implement policy iterations (Krishnamurthy et al., 2024; Brooks et al., 2023).
@@
The implementation, as depicted in Figure 3, instantiates HAI-Co2 as follows. The co-construction starts with the user providing an underspecified description of the task (in this example, building a simulation software in Python). The specification module ( 1 in Figure 3) generates the natural language abstraction of the candidate solution as a list of possible components of the software along with their functionalities  [^2] See https://subha0009.github.io/ExAIC-Interactions/SpecificationLayer.html for the details. This serves as a transparent interface in natural language that provides a layout of the construction. The user can directly edit the specification if they have specific requirements in mind (preemptive reviewer), or choose to continue with the workflow and decide on the specifics upon observing the final program (lazy reviewer).
@@
Next, the UML generation module generates a set of stochastic refinements of the natural language specification into UML descriptions ( 2 in Figure 3). The UML description of the software forces the subsequent code generation module to generate a final program that consists of multiple, independent components (in this case, Python classes) and well-defined dependencies among such components. Micro-level changes to the code (e.g., changing the design of the GUI, choice of numerical algorithms in the simulator, etc.) can be facilitated now without changing the complete codebase – a desirable property of our implementation that is closer to real-life software engineering. This addresses the challenge monolithic code LLMs face in scenarios in which persistent editing is required.
To facilitate exploration of the candidate solution space, we generate multiple UML descriptions from a given specification by setting a high decoding temperature in the UML generation module and sampling multiple responses  [^3] See https://subha0009.github.io/ExAIC-Interactions/UMLLayer.html for the 16 samples we generate in this case study. Intuitively, we seek to exploit earlier findings that a highly stochastic generation regime facilitates better novelty (Wang et al., 2023a). However, generating the Python programs from all such candidate UMLs and verifying them one by one is both computationally expensive and infeasible for the human user.
@@
To facilitate partial automation of searching through the candidate UMLs, we simulate human preference using GPT-4 Turbo in the preference learning module ( 3 in Figure 3). Specifically, the task of the preference learning module is to provide judgment on the relative quality of a given pair of UML candidates along with a justification of the preference. We simulate a tournament ( 4 in Figure 3) among the candidate UMLs by iteratively declaring one among a pair as the winner of a round. After a logarithmic order of such rounds ($\log_{2}n$ being the depth of the tournament tree for $n$ candidate UMLs), the preference learning layer comes up with a final pair and a summary of preference justifications  [^4] See https://subha0009.github.io/ExAIC-Interactions/PreferenceLayer.html for the tournament on the 16 UMLs generated in the first phase. Note that although we seek to minimize the cost of human intervention in this step by automating preference-based ranking, one can envision the human expert providing their judgement on these UMLs. In such a setup, the preference-learning module needs to explicitly adapt to such gold preference examples.

### v3

@@ 3.2 Latent utility
We assume that the construction space is equipped with a latent utility function reflecting the preferences of the expert, i.e., the quality of solutions as perceived by the expert. In general, “quality” may refer to various dimensions or criteria, and different objectives might be pursued at the same time; we formalize this with a multidimensional utility function $u({\bm{x}})=(u_{1}({\bm{x}}),\ldots,u_{K}({\bm{x}}))^{\top}$ comprised of local utility functions $u_{i}$. For example, a computer program could be rated by average runtime or memory consumption. The local utility functions can be combined into a scalar utility function $U:\,{\mathcal{X}}\rightarrow\mathbb{R}$ via a suitable aggregation operator.
Various factors influencing the quality of candidate solutions can be distinguished, notably hard and soft constraints. Hard constraints refer to (functional) properties that qualify a candidate as a valid solution. For example, a computer program should properly compile and not contain any syntax errors. Even if invalid solutions should normally be considered useless, the abstract notion of utility is flexible enough to distinguish different levels of invalidity. For example, a non-executable computer program may still have a non-zero utility if the error can easily be fixed by the expert. In any case, hard constraints will normally not identify a solution uniquely. For example, there are many computer programs that are functionally equivalent in the sense of having the same input-output behavior. Soft constraints refer to criteria that make a solution more or less desirable such as the length of a computer program and its time and memory consumption.
In general, the utility (be it in the form of the multidimensional utility function $u$ or the scalar utility function $U$) is not known to the AI, nor is the expert explicitly aware of it. Rather, this utility is latent and underlies the expert’s preference feedback. From this preference feedback, the AI can learn an approximation $\hat{U}$. The AI’s goal is then to construct a solution ${\bm{x}}^{*}$ that maximizes $\hat{U}$, or which is at least close to the maximizer, while simultaneously improving the approximation quality of $\hat{U}$. The utility $U$ also induces utilities on higher levels of abstraction. For example, one way to “lift” a utility function from level ${\mathcal{X}}_{j}$ to the more abstract level ${\mathcal{X}}_{j-1}$ is via aggregation: $U({\bm{x}}^{\prime})=\alpha(\{U({\bm{x}})\,|\,{\bm{x}}\in{\mathcal{X}}_{j},f_{j}({\bm{x}})={\bm{x}}^{\prime}\})$, where $\alpha$ is an appropriate aggregation function (Grabisch et al., 2009).
In our case study, we represent the utilities as user preferences in natural language, extracted from the interaction where both the user and the AI agent can choose between presented options, provide explicit instructions, or actively edit parts of the solution.
@@ 3.3 Interaction and preference-based search
To guide the search, human and AI can communicate via natural language; e.g., the AI agent may ask the expert for feedback or explicit advice. Alternatively, the expert may actively intervene, for example by critiquing or modifying a candidate solution. A third type of interaction, particularly important in the context of HAI-Co2, is driven through preferential feedback: By informing the AI agent about the quality of candidate solutions, the expert provides hints at presumably more promising (and, likewise, less promising) regions of the construction space, and hence suggests promising “search directions” to the AI agent. To give an example, the expert can compare two competing candidate solutions with each other (e.g., whether a modification has improved a solution or made it worse) and provide this feedback (in natural language) to the AI agent for the next iteration. The AI agent utilizes the feedback to improve its approximation $\hat{U}$ of the latent utility function, which is an important element of its informational state.
In our case study, for example, we implement a preference-based search strategy that identifies promising solutions via a tournament of pairwise comparisons. Besides, we realize a search policy that refines an existing solution guided by the expert’s preferences (see Section 5 for details).
The way in which the AI agent and the human expert cooperate with each other is defined in the form of a protocol. Among other things, the protocol clarifies the types of queries and responses on the two sides (AI agent and human expert) and the (preference) feedback that can be given by the expert.
@@
HAI-Co2 comprises a broad variety of human-AI interaction in natural language, as well as categorical choices and active modification of the candidate solution by the human expert. The search policy $\pi$ is designed to generate a (locally) optimal candidate solution based on the immediate as well as historical feedback, thereby adapting to the preference signals from the human expert user. The hierarchical abstraction of the search space facilitates a modular modification of the complex candidate solution. As we will see in the case study (Section 5), HAI-Co2 also allows for incorporating creative components into the generation of candidate solutions, for example through the injection of randomness in the heuristic search process or the refinement of abstract into more concrete solutions.
@@ 4 Challenges and practical issues
Specification of abstraction hierarchy. Two core components of HAI-Co2 are (i) the abstraction hierarchy of the construction space and (ii) the neighborhood structure that facilitates preference-based search. A synergistic implementation of (i) and (ii) poses a non-trivial research challenge. In the domain of code generation, Le et al. (2024) propose a modular code generation approach to circumvent this challenge: they generate a chain-of-thought style intermediate description of the subtasks followed by modular codes implementing each of them. Such a hierarchical generation approach can be extended to generalized solution co-construction. However, relying on purely natural language-based intermediate representations limits the utility of the hierarchy. The choice of abstraction representation is domain-specific: UML descriptions are a suitable abstraction for code generation, but not for an AI scientific assistant. The abstraction specification must adhere to the neighborhood structure on all levels of abstraction; e.g., if two points are neighbors, their abstractions must also be neighbors. This poses additional constraints on the choice of abstraction, the choice of neighborhood structure, and the choice of refinements between two given levels of abstraction. Ideally, all levels of abstraction must be suitable for human preference expression: an “unnatural” choice of abstraction can render the co-construction tiresome for the human expert, negatively affecting productivity.
Learning from active human edits. The role of the AI agent as a co-construction partner is central to HAI-Co2. This entails the possibility of active participation from the human expert and the need to learn expert preferences from such participation. Current generative AI lacks the necessary structures of the solution space, primarily represented in its input context, that could delineate the changes introduced by the expert and, subsequently, be the basis for learning from it. HAI-Co2 provides a plausible alternative to “put everything in context” that can solve this challenge, as we argue in the following. Let ${\bm{x}}\in{\mathcal{X}}_{j}$ and ${\bm{x}}^{\prime}\in{\mathcal{X}}_{j}$ be the solution before and after the edit from the human expert. The neighborhood structure imposed by HAI-Co2 on ${\mathcal{X}}_{j}$ requires learning the changes in the utility function upon moving from a candidate solution to a neighboring one. If one ensures a vector space structure formed by the utility ${\bm{u}}({\bm{x}})$, then the expert preference is equivalent to ${\bm{u}}({\bm{x}}^{\prime})-{\bm{u}}({\bm{x}})$. Even without learning to map the solutions to the utility, one can simply seek to learn the mapping from ${\bm{x}}^{\prime}-{\bm{x}}$ to expert preferences. Gao et al. (2024) previously showed that learning preferences from such changes is superior to prompting-based methods in terms of aligning LLMs to user edits. However, their experiments are focused on simpler, general-purpose natural language generation tasks. In expert-domain applications where the utility of a solution includes multiple hard constraints (e.g., executability of code) along with stylistic preferences, learning a structured representation of the utility space is essential and a challenge on its own.
@@
Specification of informational state. HAI-Co2 utilizes an informational state to keep track of relevant information in the interaction history. Given that the search policy is conditioned on it, an efficient representation of the informational state is a core challenge of HAI-Co2. Typically, such interactive co-constructions are expected to span a long sequence context. While we have observed a significant surge in the context-size of present-day generative AI (e.g., GPT-4 Turbo can handle up to 128K tokens in the input prompt), recent research has questioned the effective usability of such very long context information (Liu et al., 2024). The representation of the informational state needs to be compatible with the abstraction specification of the construction space as well as the choice of how preference signals from the human expert are encoded. This is particularly important as the reflection of any preference signal upon the candidate solution is manifested via the informational state – an unreliable update of the informational state subsequently worsens the solution quality and may result in a divergent search.
@@
Communication in natural language. When humans co-construct a solution, communication in natural language plays an important role. Natural language is a powerful and at the same time succinct medium for conveying information. Given the expressivity of natural language, human and AI agent can easily communicate different options of how to improve the current solution, both at a detailed level and in more abstract terms. Similarly, preference learning is facilitated by natural language, since many preferences are easily specified in natural language. The challenge here is that the language capabilities of LLMs have advanced to an impressive level for the general domain, but this does not apply to complex expert domains (Magesh et al., 2024; Hager et al., 2024; Anand et al., 2024).
@@
Multimodal human-AI interaction. Natural language-based interaction is not the ideal channel for all types of preferences. Categorical preference can be communicated more simply by pointing towards a better solution. Thus, we would like to incorporate multiple types of preference into HAI-Co2. Similar to deciphering the preference from natural language, different modalities and modeling approaches have their own sets of challenges and require non-trivial research efforts. For example, inferring a preference-based global ranking from pairwise comparisons can be challenging. Popular methods like the Bradley-Terry model (Bradley & Terry, 1952) have their own limitations, such as strong assumptions on the preference structure. Prior literature tackling such hurdles (Mao et al., 2018; Shah et al., 2016) paves the way for research under the umbrella of HAI-Co2. Additionally, incorporation of expert preferences across multiple modalities poses the challenge of aligning these multiple modes of feedback with each other. For example, the human expert may express the need for a security feature in a software engineering problem explicitly, or they can express it implicitly by choosing a candidate solution that includes the feature over one that does not. The AI needs to extract equivalent preference information in these two scenarios. Contemporary research in recommender systems that deal with modeling user preferences on multiple item modalities (Guo et al., 2018; Xu et al., 2021) can serve as a starting point. However, the relative complexity and nuances of preferences in the case of HAI-Co2 hinder a trivial extension of recommendation-oriented solutions.
@@
Dynamic user preferences. Current techniques of aligning neural AI systems to human preferences, broadly referred to as RLHF (Reinforcement Learning from Human Feedback), typically involve a two-stage process: learning a reward model on preference data followed by fine-tuning a foundation model (often an LLM or a diffusion model) upon reward supervision from the reward model (Kaufmann et al., 2024). This setup is fundamentally limited to static adaptation in the regime of expert-domain co-construction; a single model of human preferences is imitated by the agent that cannot adapt to the personalized preferences of the human expert. In practice, user preferences are dynamic and evolve over time (Franklin et al., 2022). They can also be influenced by and hence be constructed during the elicitation process itself (Lichtenstein & Slovic, 2006). This is a fundamental challenge in co-construction problems, where the AI agent must adapt to the evolving preferences of the human expert. Multi-turn RLHF (Zhou et al., 2024), though it extends the context of preference-adherence to an iterative, conversational regime, cannot solve the challenge of dynamically evolving user preferences. The PAL framework, proposed by Chen et al. (2024), provides a partial solution to our problem via personalized modeling of static human preferences. Unlike traditional policy learning, HAI-Co2 motivates a reward-free exploration of the solution space (Jin et al., 2020). In-context reinforcement learning can pave the way towards handling dynamic preference signals (Yang et al., 2024; Lee et al., 2024). However, the action space in the scope of HAI-Co2 overlaps with the generation of multiple hierarchical views of the candidate solution, rendering the problem much harder than existing work on in-context RL. Prior work with LLMs showcases the possibilities of using them as in-context agents, though exploration abilities will need fine-tuning-based interventions (Krishnamurthy et al., 2024).
@@
  • Quality of the solution should be evaluated using domain-specific measures; irrespective of the process of co-construction or human preferences, the solution must fulfill some objective criteria of correctness.
@@
  • Complexity of co-construction includes the computational complexity of generation and preference-based search – resulting in potentially high computational cost – and the cognitive complexity of the framework – resulting in cognitive load for the expert user. The latter demands significant research efforts from a multidisciplinary approach to ensure that automated assistants truly bring value to the expert.
@@
Given that human experts are costly and have limited time, LLM-based simulation of human-AI interaction may facilitate large-scale evaluation (Tamoyan et al., 2024). Even though our four evaluation criteria seem to demand human evaluation, we conjecture that the development of artificial critic models (McAleese et al., 2024), with human value alignment, will be an important research direction in the future.
Figure 3: Instantiation of HAI-Co2 for the problem of building a double pendulum simulation. A solution is co-constructed through human-AI cooperation as follows. The interaction between the human and the AI is shown in the red box on the right, top to bottom. We define the co-construction space on three levels of hierarchy: Specification space, UML space, and Python program space. The user starts by specifying the (potentially incomplete) problem to solve. The Specification module (green rectangle) generates a pair of candidate specifications of the software to build (S1 and S2) and presents them to the user. The user expresses their preference in two manners: i) they choose one of the candidate solutions (S1) as better than the other, and ii) provide partial edits to the specification directly. The Preference decoder module (purple rectangle) extracts preference rules from the interaction context. Based on the decoded preferences, the Specification module generates a new pair of candidate specifications (only one shown for space reasons), from which the user chooses one. The UML generation module serves as a generator of refinements from specification space to UML space and generates a set of four UMLs from the specification selected. The Preference-based search module then runs a tournament-based search among these UML candidates: a pair of UMLs are compared against the specification and the decoded user preferences and one is chosen. Two finalist UMLs from the tournament are then used by the Code generation module (pink rectangle) to generate two candidate Python programs. These programs are presented to the user again for their feedback.
5 An exemplary simulation of HAI-Co2
@@
In this section, we provide a case study of HAI-Co2, tailored to code generation as a co-construction problem. This case study does not claim scientific rigor on its own; instead, we use it to demonstrate what prior findings (see Section 1 and Section 2) already establish. We do not provide a complete implementation of HAI-Co2; in particular, the following are not included in the case study: neighborhood structure of the solution space, preference extraction from actual human participation, dedicated utility function tailored toward the expert problem. Instead, we emulate the intended behavior of a complete implementation using prompted LLMs, with the goal of motivating the practicality of HAI-Co2.
@@
The initial problem description is underspecified. During the cooperation, the user can introduce new requirements, ask for modifications to the already generated code, and so on. We approximate different aspects of HAI-Co2 (the surjective mappings between different abstraction hierarchies of the construction space, policy and heuristic search strategy) using baseline implementation strategies for ease of demonstration. Future research endeavors should be directed to more in-depth implementation of these features.
@@
In this example, the construction space ${\mathcal{X}}$ consists of the set of all Python programs. Three distinct levels of abstraction are implemented. (i) Specification space. A specification of the simulation software in natural language (${\mathcal{X}}_{0}$). (ii) UML space. A UML description of the software (${\mathcal{X}}_{1}$). (iii) Python program space. The Python program (${\mathcal{X}}_{2}$) itself. The abstraction refinements $f^{(-1)}_{1}$ and $f^{(-1)}_{2}$ (as introduced in Section 3) are implemented using suitably prompted instances of GPT-4 Turbo that we denote as UML generation module and Code generation module, respectively; while the former produces a (stochastic) set of refinements in UML given a natural language specification, the latter generates Python implementations of a given UML description. To decode the user’s preferences from the interaction, we use a Preference decoder module, implemented using prompted GPT-4 Turbo. Following the focus on natural language, the informational state ${\mathcal{I}}$ is realized primarily as the interaction history in natural language, along with an explicit list of preference rules decoded from this interaction. One can impose a geometric structure on ${\mathcal{I}}$ by introducing explicit metric space structure on the different abstraction spaces (e.g., edit distance), rendering ${\mathcal{I}}$ to behave like a trajectory. However, introduction of such structures will be dependent on the expert domain application.
To facilitate the exploration of the candidate solution space, we generate multiple solution representations on different abstraction levels by setting a high decoding temperature in the respective generation modules and sampling multiple responses. Intuitively, we seek to exploit earlier findings that a highly stochastic generation regime facilitates better novelty (Wang et al., 2023a). Furthermore, this imposed stochasticity can be interpreted as the notion of neighborhood in the respective spaces: one can treat two solutions sampled from the same input context of the generation module as neighbors; distance between two different input contexts can be measured by edit distance. Although we do not explicitly specify such geometric structure in this example, the search strategy uses it implicitly.
We do not implement a concrete realization of the search policy $\pi$; instead, we rely on the limited abilities of LLM instances to explore and implement policy iterations (Krishnamurthy et al., 2024; Brooks et al., 2023). While the notion of search is present across all three levels of abstraction, we perform explicit search in the UML space using the Preference-based search module, which runs a tournament among candidate UML solutions, guided by the decoded preferences.
@@
The implementation, as depicted in Figure 3, instantiates HAI-Co2 as follows. The co-construction starts with the user providing an underspecified description of the task (in this example, building a simulation software in Python). The specification module (green rectangle in Figure 3) generates the natural language abstraction of the candidate solution as a list of possible components of the software along with their functionalities. This serves as a transparent interface in natural language that provides a layout of the construction. A pair of candidate specifications are generated using a high-temperature stochastic generation regime. The user chooses one of them as better. Additionally, they can state any explicit modification request. Moreover, they can directly edit the specification if they have specific requirements in mind (preemptive reviewer), or choose to continue with the workflow and decide on the specifics upon observing the final program (lazy reviewer). The Preference decoder module (purple rectangle in Figure 3) lists down the preferences decoded from the user’s actions. If the user introduces any new modification (e.g., in the presented example, they remove certain modules from the specification and insert new modules), the specification module generates a new pair of specifications for the user to provide feedback on. This continues till a suitable specification is obtained.
@@
Next, the UML generation module generates a set of stochastic refinements of the natural language specification into UML descriptions. The UML description of the software forces the subsequent code generation module to generate a final program that consists of multiple, independent components (in this case, Python classes) and well-defined dependencies among such components. Micro-level changes to the code (e.g., changing the design of the GUI, choice of numerical algorithms in the simulator, etc.) can be facilitated now without changing the complete codebase – a desirable property of our implementation that is closer to real-life software engineering. This addresses the challenge monolithic code LLMs face in scenarios in which persistent editing is required. However, generating the Python programs from all such candidate UMLs and verifying them one by one is both computationally expensive and infeasible for the human user.
@@
The preference-based search among the candidate UMLs is implemented as a tournament by iteratively declaring one among a pair as the winner of a round. After a logarithmic order of such rounds ($\log_{2}n$ being the depth of the tournament tree for $n$ candidate UMLs), the Preference-based search module comes up with a final pair and a summary of preference justifications. [^6] See https://subha0009.github.io/ExAIC-Interactions-10A2/PreferenceLayer.html for the tournament on the candidate UMLs generated Note that although we seek to minimize the cost of human intervention in this step by automating preference-based ranking, one can envision the human expert providing their judgement on these UMLs. In such a setup, the Preference-decoding module can be used to explicitly adapt to such gold preference examples.